INTERMOLECULAR POTENTIAL AND ROTATIONAL SPECTRA OF H_{2}-OCS COMPLEXES

\underline{KELLY J. HIGGINS}, ZHENHONG YU, and WILLIAM KLEMPERER, Department of Chemistry and Chemical Biology, Harvard University, Cambridge, Massachusetts 02138; MICHAEl C. Mc-CARTHY and PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics, Cambridge, Massachusetts 02138 and Division of Engineering and Applied Sciences, Harvard University, Cambridge, Massachusetts 02138; KRISTINE LIAO and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB, Canada T6G 2G2.

The interaction between molecular hydrogen and carbonyl sulfide was studied through ab initio calculations and microwave spectroscopy of pH_{2}-OCS, oH_{2}-OCS, pD_{2}-OCS, oD_{2}-OCS, and HD-OCS. The intermolecular potential surface (IPS) encompasses all four intermolecular degrees of freedom and is an extensive refinement of the IPS previously presented.^{a} The IPS was calculated at the MP4/aug-cc-pVTZ + bond functions level of theory at a total of 1836 unique geometries. The global minimum is -210.3 cm^{−1}and places the hydrogen on the side of the OCS in a near parallel arrangement. The interaction is dominated by dispersion with little contribution from the electrostatic dipole-quadrupole term. Four dimensional bound state calculations using this IPS yield binding energies of -76.7 cm^{−1}for pH_{2}-OCS and -90.3 cm^{−1}for oH_{2}-OCS relative to j = 0 or 1 H_{2}, respectively. Surprisingly, the ground state for all species, including oH_{2}-OCS and pD_{2}-OCS with hydrogen angular momentum j = 1, have total angular momentum J = 0.

Eleven to fifteen a- and b-type pure rotational transitions were measured for each of the five species using a Fourier transform microwave spectrometer. Careful control of the gas mixture was required to observe pH_{2}-OCS in the presence of the more strongly bound oH_{2}-OCS species. The observed transition frequencies of each species can be fit using a standard asymmetric rotor Hamiltonian with the exception of oH_{2}-OCS, for which the effects of internal rotation require a more complicated treatment. Comparison will be made between the ab initio calculated and the observed transition frequencies and fitted spectroscopic constants.

aHiggins, Yu, and Klemperer 57th International Symposium on Molecular Spectroscopy (2002)